OpenStudy (wyattp17):

a 24 foot long bleacher stand has a base angle of 30 degrees how high above the ground is the last row of seating

OpenStudy (amonoconnor):

SIN(theta) = y/r= opposite or, the height of the triangle/hypotenuse So... SIN(30) = y/(24) (.5773502692) = y/(24) y = (.55735...)24 = 12 feet